OpenStudy (anonymous):

Find the particular solution y=f(x) to the given differential equation with the initial condition f(0)=-2 dy/dx=e^x+4x/y

OpenStudy (anonymous):

Is the right side \(\dfrac{e^x+4x}{y}\), or \(e^x+\dfrac{4x}{y}\) ?

OpenStudy (anonymous):

\[\frac{ dy }{ dx}=\frac{ e^x+4 }{y }\]

OpenStudy (anonymous):

Oh, much easier. It's separable, so you can split the differentials: \[y~dy=(e^x+4)~dx\]

OpenStudy (anonymous):

Integrate both sides, then make use of the initial condition. Solve for \(y\) if you like

OpenStudy (mathmale):

Great work, Siths! But please involve @sarcos11 towards finding his/her own solution. I'm sure he/she can do some of this work unaided. Thank you.

OpenStudy (anonymous):

No problem! I have trouble staying my hand most of the time :)

OpenStudy (anonymous):

I'm a girl.

OpenStudy (anonymous):

my teacher taught me that we are suppose to use u ex. u=y du/dy=1 dy=du/1 is this necessary for this problem

OpenStudy (unklerhaukus):

do you mean \[\frac{ dy }{ dx}=\frac{ e^x+4x }{y }\]?

OpenStudy (mathmale):

You asked, " u=y du/dy=1 dy=du/1 is this necessary for this problem" I'd agree with sarcos11: no. \[\frac{ dy }{ dx}=\frac{ e^x+4 }{y }\] is very easily separable, as sarcos11 has already pointed out. Separate the equation by variables, and then integrate each side separately.

OpenStudy (anonymous):

i'm sorry i meant to say yes

OpenStudy (anonymous):

I am so confused right now

OpenStudy (anonymous):